OpenStudy (anonymous):

Joan and amber have a combined age of 34. If amber is 2 years less than twice the age of joins age, how old is each person?

OpenStudy (anonymous):

put m as amber's age, m as amber's. then \[m+n=34\] \[m=2n-2\] \[2n-2+n=34\] solve for n

OpenStudy (anonymous):

3n - 2 = 34

myininaya (myininaya):

so i guess n is joan's age lol